# Spectral Bottleneck in Deep Neural Networks: Noise is All You Need

### v1

@@ Abstract
Deep neural networks are known to exhibit a spectral learning bias, wherein low-frequency components are learned early in training, while high-frequency modes emerge more gradually in later epochs. However, when the target signal lacks low-frequency components and is dominated by broadband high frequencies, training suffers from a spectral bottleneck, and the model fails to reconstruct the entire signal, including the frequency components that lie within the network’s representational capacity. We examine such a scenario in the context of implicit neural representations (INRs) with sinusoidal representation networks (SIRENs), focusing on the challenge of fitting high-frequency-dominant signals that are susceptible to spectral bottleneck. To effectively fit any target signal irrespecitve of it’s frequency content, we propose a generalized target-aware weight perturbation scheme (WINNER - weight initialization with noise for neural representations) for network initialization. The scheme perturbs uniformly initialized weights with Gaussian noise, where the noise scales are adaptively determined by the spectral centroid of the target signal. We show that the noise scales can provide control over the spectra of network activations and the eigenbasis of the empirical neural tangent kernel. This method not only addresses the spectral bottleneck but also yields faster convergence and with improved representation accuracy, outperforming state-of-the-art approaches in audio fitting and achieving notable gains in image fitting and denoising tasks. Beyond signal reconstruction, our approach opens new directions for adaptive weight initialization strategies in computer vision and scientific machine learning.
1 Introduction
@@
Figure 1: Spectral bottleneck issue in SIREN and overview of the weight perturbation scheme. a, We attempt to fit eight discretely sampled broadband 1D signals (S1–S8) with decreasing low-frequency content. As shown in the right panel, the PSNR of SIREN sitzmann2020implicit progressively decreases from S1 to S4, eventually encountering a spectral bottleneck after S4. SIREN fails to capture nearly all the frequencies of signals S4–S8, even though its frequency support should in principle allow it to represent a portion of the spectrum. In contrast, SIREN2 initialized with WINNER maintains higher PSNR across all signals. b, Schematic of a feedforward neural network with periodic activations, illustrating the statistical distributions of layer-1 outputs $X_{ij}$, weight matrix $W_{jk}$, and noise matrix $\eta_{jk}$. The effect of Gaussian noise ($\eta_{jk}$) on the spectrum of layer 2 pre-activations is shown: WINNER enhances the receptivity of high-frequencies.
@@
Implicit neural representations (INRs) of natural and synthetic signals have broad applications across various domains. They allow neural networks to represent coordinate-based discrete data such as images, videos, audio, 3D shapes, and scientific datasets as continuous functions. This allows seamless integration of scientific, multimedia, and medical data into machine learning pipelines for tasks such as denoising, classification, inpainting, and latent representation mildenhall2021nerf ; oechsle2019texture ; niemeyer2020differentiable ; szatkowski2022hypersound . The key advantages of INRs over discretely sampled data are continuous input parameterization (no grid), fully differentiable networks with accessible gradients, and compressed data representation sitzmann2020implicit ; tancik2020fourier . With spatio-temporal derivatives readily available through automatic differentiation, INRs can also be employed to solve forward and inverse problems governed by differential equations in a mesh-free setting raissi2019physics ; song2022versatile .
@@
Although training a conventional multi-layer perceptron (MLP) to learn discrete data through supervised learning may seem straightforward, representing natural signals (such as images, videos, audio, or turbulent fluid flow data) is often challenging due to the wide range of frequencies and rank complexity of the data. Several works have theoretically and empirically showed that ReLU based deep neural networks are prone to ‘spectral bias’ rahaman2019spectral ; xu2019frequency and are ill-conditioned for low-dimensional coordinate-based training tasks. To overcome spectral bias and the associated lazy training, coordinate inputs are often mapped to positional encodings mildenhall2021nerf or random Fourier features rahimi2007random ; tancik2020fourier , which enable more effective learning of high-frequency signals. Alternatively, Sitzmann et al. sitzmann2020implicit have proposed sinusoidal representation networks (SIRENs) using periodic activation function $\phi(x)=\sin(\omega_{0}x)$,
$f^{\text{SIREN}}(\mathbf{x};\theta)=\mathbf{W}^{(L)}\mathbf{h}^{(L-1)}+\mathbf{b}^{(L)},\quad\mathbf{h}^{(l)}=\phi^{\text{sin}}\left(\mathbf{W}^{(l)}\mathbf{h}^{(l-1)}+\mathbf{b}^{(l)}\right),\quad\mathbf{h}^{(0)}=\mathbf{x},$ (1)
@@
where $\mathbf{x}\in\mathbb{R}^{d}$ is the input, $\theta=\{\mathbf{W}^{(l)},\mathbf{b}^{(l)}\}_{l=1}^{L}$ are the network parameters, L1L-1 hidden layers, and $\omega_{0}$ is the activation periodicity specified as hyperparameter. They also propose a principled initialization scheme for the weights $\mathbf{W}$, to ensure network’s pre- and post-activations at initialization remain narrowly bounded. Their approach faithfully captures both the discrete data and its gradients with high-fidelity, enabling applications in computer vision and the solution of scientific differential equations. Various other variants related to SIREN were also proposed, such as ziyin2020neural ; liu2013fourier ; silvescu1999fourier . SIRENs were also demonstrated to significantly mitigate the spectral bias in Physics Informed Neural Network (PINN) related applications for solving initial and boundary value problems governed by ordinary and partial differential equations wang2022and ; wang2023expert ; song2022versatile ; aghaee2024performance ; raissi2019physics .
@@
While SIRENs with the principled uniform weight initialization scheme sitzmann2020implicit perform well for fitting images and videos, they struggle with low-dimensional signals such as audio when the contribution of low frequencies in the target signal is small. The reconstruction accuracy is closely tied to the signal’s spectral content; for example, under the same initialization, SIRENs have difficulty fitting signals dominated by either very high or very low frequencies. Prior works such as mehta2021modulated ; liu2024finer also suggests that SIRENs suffer with overfitting issues as the signal length increases (or in other words the contribution of high-frequencies increases). An easy way to get around this problem is to increase the input dimensionality by mapping the input coordinates to a random Fourier feature space rahimi2007random ; tancik2020fourier ; mildenhall2021nerf . However, positional embeddings lead to a quadratic increase in the parameter count with respect to the embedding dimension, and consequently with the hidden-layer width. We propose a noisy weight initialization scheme that guides the training dynamics to prevent the SIREN architecture from failing to represent frequency components of a given target. Figure 1 illustrates the outline of our noise initialization scheme and the spectral bottleneck issue of standard SIREN sitzmann2020implicit , comparing their representation accuracy to that of our proposed method on several synthetically generated audio signals, each containing 150,000 samples. All experiments in Figure 1 use an MLP with four hidden layers with 222 features in each layer.
Related work. Mathematical analysis of Basri et al. basri2020frequency ; ronen2019convergence have showed that for 1D signals, the convergence rate under gradient descent scales inversely with the square of the target signal’s frequency (i.e., $1/k^{2}$), and this frequency-dependent slowdown grows exponentially with the input dimension bietti2019inductive ; cao2019towards . More recent works like FINER liu2024finer and FINER++ zhu2024finerplusplus explore bias initialization strategies to reduce the eigenvalue decay in the empirical NTK, thus increasing effective frequency support. The works of Tang et al. tang2025structured and Varre et al. varre2023spectral (for linear networks) demonstrate how network initialization can affect parameter optimization. Building on these insights, a variety of models have been proposed to enhance the ability of neural networks to represent audio signals choudhury2024nerva ; kim2023regression ; kim2023generalizable ; li2024asmr . Techniques such as hypernetworks, positional encoding, and auxiliary networks have also been explored to improve reconstruction fidelity and reduce reconstruction noise marszalek2025hypernetwork ; lanzendorfer2023siamese . Audio-specific works include Siamese SIREN lanzendorfer2023siamese for compression, HyperSound szatkowski2022hypersound for INR generation via meta-learning, and INRAS su2022inras for spatial audio modeling. Neural audio representations continue to find applications in classification, speech synthesis, sound event detection, encoding, and embedding sharan2021benchmarking ; tessarini2022audio ; hershey2017cnn ; donahue2018adversarial ; shen2018natural ; aironi2021graph .
@@
2 Understanding ‘Spectral Bottleneck’
2.1 Challenges of Fitting 1D Signals with SIREN
@@
As for SIREN sitzmann2020implicit , increasing the activation periodicity $\omega_{0}$ (Eqn 1) may seem like a potential solution to mitigate the frequency bias induced by input scaling. However, since the pre-activation values of a SIREN predominantly lie within the range of [3,3][-3,3] due to it’s weight initialization scheme, a high $\omega_{0}$ can render the activation function insensitive and alter the activation distribution. To address this, Sitzmann et al. sitzmann2020implicit scaled the inputs by a factor of 100, sampling them from a wider range $\mathbf{x}\sim\mathcal{U}(-100,100)$, which effectively increases the power spectral density of the pre-activations by approximately the same factor (see Proposition 1). This strategy, while effective for low-frequency-dominant signals, we observe SIREN with input scaling still fails for signals dominated by high frequencies. An example illustrating this failure is discussed in Fig. 2.
@@
An example case when SIREN fails to fit a high-frequency signal. We consider the reconstruction of a high-frequency audio signal $f:\mathbb{R}^{2}\to\mathbb{R}^{d}$, using the example tetris.wav available in the linked GitHub repository. Experiments conducted using SIREN in Fig. 2(a,b) show that, although input scaling improves PSNR value, significant errors remain in the reconstructed signal, as shown in the spectrogram error. A maximum PSNR value of just $\sim$23.5dB was achieved using a scaling factor of $3\times 10^{4}$. We also experimented by increasing network size (Fig. 2b), changing scheduler parameters, and changing the frequency parameter $\omega_{0}$, none of which resolve this issue. This reveals that SIREN, with its default weight initialization, struggles to fit such high-frequency dominant signals due to the spectral bottleneck phenomenon. Other 1D examples are shown in Fig. 1b, where SIREN exhibits a spectral bottleneck for signals S4–S8, all lacking low-frequency components.
Remark.  Repeating the experiment in Fig. 2 with alternative architectures, including WIRE saragadam2023wire , FINER liu2024finer , HOSC serrano2024hosc , and Gauss ramasinghe2022beyond , all yield the same outcome: convergence to a fixed PSNR of 13.4 with a zero-valued output. This pathology is therefore not just confined to SIRENs, but is shared across other related deep neural networks. We further observed that mapping the inputs to random Fourier features rahimi2007random ; tancik2020fourier or adopting a broader bias initialization, as proposed in zhu2024finerplusplus , mitigates this failure mode. A formal characterization of this behavior is presented in Sec. 2.2.
2.2 Learning Dynamics in Fourier Space
@@
Figure 4: Contrasting learning dynamics of SIREN for low and high frequency-dominant targets. a Output evolution when fitting the low-frequency signal of Eqn. 5. The right subplot shows the NTK spectral energy, $\log(1+\mathcal{S}(k))$, across frequency eigenmodes $k$ during training, exhibiting a steady increase that indicates effective learning. b Output and NTK spectral energy evolution for a high-frequency signal. The left subplot shows that SIREN fails to match the ground truth, while the right subplot reveals a suppression of spectral energy over training, indicating difficulty in representing high-frequency-dominant signals.
@@
As shown in Fig. 4, SIREN accurately reconstructs the low-frequency target, with $\mathcal{S}(k)$ progressively increasing in the relevant modes. However, for the high-frequency target, the output remains near zero and $\mathcal{S}(k)$ is steadily suppressed, indicating an inability to fit high-frequency components. To address this phenomenon, we introduce a weight perturbation scheme in section 3.
@@
Remark.  A key aspect of the spectral bottleneck is that the network fails to capture even the low-frequency components (circled in Fig. 3), despite these frequencies lying well within the nominal support of SIREN.
3 WINNER: Weight Initialization with Noise for NEural Representations

### v2

# Spectral Bottleneck in Sinusoidal Representation Networks: Noise is All You Need

@@ Abstract
This work identifies and attempts to address a fundamental limitation of implicit neural representations with sinusoidal activation. The fitting error of SIRENs is highly sensitive to the target frequency content and to the choice of initialization. In extreme cases, this sensitivity leads to a spectral bottleneck that can result in a zero-valued output. This phenomenon is characterized by analyzing the evolution of activation spectra and the empirical neural tangent kernel (NTK) during the training process. An unfavorable distribution of energy across frequency modes was noted to give rise to this failure mode. Furthermore, the effect of Gaussian perturbations applied to the baseline uniformly initialized weights is examined, showing how these perturbations influence activation spectra and the NTK eigenbasis of SIREN. Overall, initialization emerges as a central factor governing the evolution of SIRENs, indicating the need for adaptive, target-aware strategies as the target length increases and fine-scale detail becomes essential. The proposed weight initialization scheme (WINNER) represents a simple ad hoc step in this direction and demonstrates that fitting accuracy can be significantly improved by modifying the spectral profile of network activations through a target-aware initialization. The approach achieves state-of-the-art performance on audio fitting tasks and yields notable improvements in image fitting tasks.
I Introduction
@@
Figure 1: Spectral bottleneck issue shown using SIREN and overview of the weight perturbation scheme. (a) We attempt to fit eight discretely sampled broadband 1D signals (S1–S8) with decreasing low-frequency content. As shown in the right panel, the PSNR of SIREN sitzmann2020implicit progressively decreases from S1 to S4, eventually encountering a spectral bottleneck after S4. SIREN fails to capture nearly all the frequencies of signals S4–S8, even though its frequency support should in principle allow it to represent a portion of the spectrum. In contrast, SIREN2 initialized with WINNER maintains higher PSNR across all signals. (b) Schematic of a feedforward neural network with periodic activations, illustrating the statistical distributions of layer-1 outputs $X_{ij}$, weight matrix $W_{jk}$, and noise matrix $\eta_{jk}$. The effect of Gaussian noise ($\eta_{jk}$) on the spectrum of layer 2 pre-activations is shown: WINNER enhances the receptivity of high-frequencies.
@@
Implicit neural representations (INRs) of natural and synthetic signals have broad applications across various domains. They allow neural networks to represent coordinate-based discrete data such as images, videos, audio, 3D shapes, and scientific datasets as continuous functions. This allows seamless integration of multimedia, medical, and scientific data into machine learning pipelines for tasks such as denoising, classification, inpainting, and latent representation mildenhall2021nerf; oechsle2019texture; niemeyer2020differentiable; szatkowski2022hypersound. The key advantages of INRs over discretely sampled data are continuous input parameterization (no grid), fully differentiable networks with accessible gradients, and compressed data representation sitzmann2020implicit; tancik2020fourier. Using the readily available spatio-temporal gradients obtained through automatic differentiation, INRs can also be employed to solve forward and inverse problems governed by differential equations in a mesh-free setting raissi2019physics; song2022versatile.
@@
Although multi-layer perceptrons (MLPs) are universal function approximators, training them to fit natural signals such as images, videos, or audio is often challenging. Analytical and empirical results are presented in previous studies rahaman2019spectral; xu2019frequency showing that ReLU based deep neural networks are prone to ‘spectral bias’ and are ill-conditioned for low-dimensional coordinate-based training tasks. To overcome spectral bias and the associated lazy training, coordinate inputs are often mapped to positional encodings mildenhall2021nerf or random Fourier features rahimi2007random; tancik2020fourier, which as shown in yuce2022structured increase the ‘expressive power’ of INRs. Alternatively, Sitzmann et al. sitzmann2020implicit have proposed sinusoidal representation networks (SIRENs) using periodic activation function $\phi(x)=\sin(\omega_{0}x)$,
$f^{\text{SIREN}}(\mathbf{x};\theta)$ $=\mathbf{W}^{(L)}\mathbf{h}^{(L-1)}+\mathbf{b}^{(L)},$ (1)
$\mathbf{h}^{(l)}$ $=\phi^{\text{sin}}\!\left(\mathbf{W}^{(l)}\mathbf{h}^{(l-1)}+\mathbf{b}^{(l)}\right),\quad\mathbf{h}^{(0)}=\mathbf{x}.$
@@
where $\mathbf{x}\in\mathbb{R}^{d}$ is the input, $\theta=\{\mathbf{W}^{(l)},\mathbf{b}^{(l)}\}_{l=1}^{L}$ are the network parameters, L1L-1 hidden layers, and $\omega_{0}$ is the activation periodicity specified as hyperparameter. They also propose a principled initialization scheme for the weights $\mathbf{W}$, to ensure network’s pre- and post-activations at initialization remain narrowly bounded. Their approach faithfully captures both the discrete data and its gradients with high-fidelity, enabling applications in computer vision and the solution of scientific differential equations. Various other variants related to SIREN were also proposed, such as ziyin2020neural; liu2013fourier; silvescu1999fourier.
@@
While SIRENs with the uniform weight initialization scheme sitzmann2020implicit can mitigate spectral bias and effectively fit 2D images, videos, and 3D geometries, they often perform poorly on low-dimensional signals such as audio when low-frequency components contribute minimally to the target. As pointed out in yuce2022structured, the capacity of network to represent a target signal’s frequency components does not guarantee efficient learning of that signal. The reconstruction accuracy is rather closely tied to the spectral content of the signal and the inductive bias of network initialization; for instance, under fixed initialization, SIRENs struggle with signals dominated by very high or very low frequencies. Prior work mehta2021modulated; liu2024finer shows that SIRENs exhibit overfitting as signal length and high-frequency content increase. An easy way to get around this problem is to increase the input dimensionality by mapping the input coordinates to a random Fourier feature space rahimi2007random; tancik2020fourier; mildenhall2021nerf. However, positional embeddings lead to a quadratic increase in the parameter count with respect to the embedding dimension, and consequently with the hidden-layer width. In this work, we examine the fitting challenges of SIREN on increasingly high-frequency targets and characterize the limitations of its traditional weight initialization sitzmann2020implicit, particularly the spectral bottleneck phenomenon, where SIREN fails to recover all frequency modes and collapses to a zero-valued function. We then propose a noisy weight initialization scheme WINNER that aims to mitigate the inductive bias at initialization. Figure 1 illustrates the spectral bottleneck issue of the standard SIREN sitzmann2020implicit and the proposed weight initialization scheme which is being employed in SIREN2. Representation accuracy of SIREN is compared with the proposed method on several synthetically generated audio signals, each containing 150,000 samples. Although, the spectral bottleneck phenomenon in Fig. 1 is demonstrated through synthetic signals, such a phenomenon was noted to also observe in fitting natural signals, examples include tap.wav and relay.wav audio clips which can be accessed through the project’s repository (link in first page). All experiments in Figure 1 use a SIREN architecture with four hidden layers with 222 features in each layer.
@@
II Related work
Activation functions for INRs. Sinusoidal activation functions have been considered in neural networks since 1999 sopena1999neural, but such models were often regarded as difficult to train parascandolo2016taming. Sitzmann et al. sitzmann2020implicit addressed this by introducing SIREN, which uses a tailored initialization scheme that enhances stable training of sinusoidal INRs. Later methods designed alternative periodic or adaptive activations jagtap2022deep to extend frequency support, such as FINER liu2024finer, FINER++ zhu2024finerplusplus, and HOSC serrano2024hosc, while others explored Gaussian or kernel-inspired nonlinearities ramasinghe2022beyond; chng2022gaussian. DINER zhu2024disorder further demonstrated that near-lossless signal representation can be achieved by augmenting a hash-table to a traditional INR backbone albeit with an additional memory overhead to store the hash data. Additional activation designs include Snake, a periodic trainable activation for learning periodic targets liu2020neural; WIRE, which uses Gabor wavelet activations to capture frequency and spatial locality saragadam2023wire; a sampling-theoretic case for using $\mathrm{sinc}$ activations in INRs saratchandran2024sampling; and STAF, a family of trainable sinusoidal activations morsali2025staf.
Weight/bias initialization. Recent works like FINER liu2024finer and FINER++ zhu2024finerplusplus explore bias initialization strategies to reduce eigenvalue decay in the empirical NTK, thus increasing effective frequency support. In geometry-focused INRs, SAL initializes network weights so that the signed distance function (SDF) starts as a 3D sphere atzmon2020sal; DiGS later introduced a multifrequency geometric initialization (MFGI) that extends this idea to sine activations and preserves high-frequency content ben2022digs; and in sinusoidal INRs, SIREN showed that without a tailored scheme, sine activations fail to reconstruct signals, motivating an initialization that stabilizes optimization sitzmann2020implicit. The works of Tang et al. tang2025structured and Varre et al. varre2023spectral (for linear networks) demonstrate how initialization can govern parameter optimization. Complementary strategies have been developed for sinusoidal and general INRs: VI3NR derives an activation-agnostic, variance-preserving rule for INR regression hewa2025vi3nr; TUNER introduces frequency-sampling input initialization together with amplitude-bounded hidden-layer weights to control the model bandlimit during training novello2025tuning; and Yeom et al. yeom2024fast propose an initialization scaling for sinusoidal representation networks to speed up optimization and alleviate spectral bias. Prior designs are often specialized (geometry for SAL/DiGS and sine activations for SIREN) and typically do not target the joint distribution of pre- and post-activations across layers. In this work, we introduce a target aware weight initialization scheme that takes into account the spectral characteristics of the target as well as preserves the layer distributions for stability and accuracy across INR tasks.
Audio INRs. A variety of models have been proposed to enhance the ability of neural networks to represent audio signals choudhury2024nerva; kim2023regression; kim2023generalizable; li2024asmr. Techniques such as hypernetworks, positional encoding, and auxiliary networks have also been explored to improve reconstruction fidelity and reduce reconstruction noise marszalek2025hypernetwork; lanzendorfer2023siamese. Audio-specific works include Siamese SIREN lanzendorfer2023siamese for compression, HyperSound szatkowski2022hypersound for INR generation via meta-learning, and INRAS su2022inras for spatial audio modeling. Neural audio representations continue to find applications in classification, speech synthesis, sound event detection, encoding, and embedding sharan2021benchmarking; tessarini2022audio; hershey2017cnn; donahue2018adversarial; shen2018natural; aironi2021graph.
In addition, analytical and empirical analysis of Basri et al. basri2020frequency; ronen2019convergence indicate that for 1D signals, the convergence rate under gradient descent scales inversely with the square of the target signal’s frequency (i.e., $1/k^{2}$), and this frequency-dependent slowdown grows exponentially with the input dimension bietti2019inductive; cao2019towards. This frequency-dependent limitation has motivated several INR-specific methods. For example, FR-INR reparameterizes MLP weights as linear combinations of fixed Fourier bases to mitigate low-frequency bias, while IGA estimates a gradient transform from a sampled empirical NTK and applies it during updates to accelerate learning of high-frequency components shi2024frinr; shi2024iga.
III Understanding ‘Spectral Bottleneck’
III-A Challenges of Fitting 1D Signals with SIREN
@@
As for SIREN sitzmann2020implicit, increasing the activation periodicity $\omega_{0}$ (Eqn 1) may seem like a potential solution to mitigate the frequency bias induced by input scaling. However, since the pre-activation values of a SIREN predominantly lie within the range of [3,3][-3,3] due to its weight initialization scheme, a high $\omega_{0}$ can render the activation function insensitive and alter the activation distribution. To address this, Sitzmann et al. sitzmann2020implicit scaled the inputs by a factor of 100, sampling them from a wider range $\mathbf{x}\sim\mathcal{U}(-100,100)$, which effectively increases the power spectral density of the pre-activations by approximately the same factor (see Proposition 1). While this strategy is effective for low-frequency-dominant signals, SIREN with input scaling can still fail on high-frequency-dominant signals. An example illustrating this failure is discussed in Fig. 2.
@@
An example case when SIREN fails to fit a high-frequency signal: We consider the reconstruction of a high-frequency audio signal $f:\mathbb{R}^{2}\to\mathbb{R}^{d}$, using the example tetris.wav available in the linked GitHub repository. Experiments conducted using SIREN in Fig. 2(a,b) show that, although input scaling improves PSNR value, significant errors remain in the reconstructed signal, as shown in the spectrogram error. A maximum PSNR value of just $\sim$23.5dB was achieved using a scaling factor of $3\times 10^{4}$. We also experimented by increasing network size (Fig. 2b), changing scheduler parameters, and changing the frequency parameter $\omega_{0}$, none of which resolve this issue. This reveals that SIREN, with its default weight initialization, struggles to fit such high-frequency dominant signals due to the spectral bottleneck phenomenon. Other 1D examples are shown in Fig. 1b, where SIREN exhibits a spectral bottleneck for signals S4–S8, all lacking low-frequency components.
Remark. Repeating the experiment in Fig. 2 with alternative architectures, including WIRE saragadam2023wire, FINER liu2024finer, HOSC serrano2024hosc, and Gauss ramasinghe2022beyond, all yield the same outcome: convergence to a fixed PSNR of 13.4 with a zero-valued output. This pathology is therefore not just confined to SIRENs, but is shared across other related deep neural networks. We further observed that mapping the inputs to random Fourier features rahimi2007random; tancik2020fourier or adopting a broader bias initialization, as proposed in zhu2024finerplusplus, mitigates this failure mode. A formal characterization of this behavior is presented in Sec. III-B.
III-B Learning Dynamics in Fourier Space
@@
Figure 3: Contrasting learning dynamics of SIREN for low and high frequency-dominant targets. (a) Output evolution when fitting the low-frequency signal of Eqn. 5. The right subplot shows the NTK spectral energy, $\log(1+\mathcal{S}(k))$, across frequency eigenmodes $k$ during training, exhibiting a steady increase that indicates effective learning. (b) Output and NTK spectral energy evolution for a high-frequency signal. The left subplot shows that SIREN fails to match the ground truth, while the right subplot reveals a suppression of spectral energy over training, indicating difficulty in representing high-frequency-dominant signals.
@@
As shown in Fig. 3, SIREN accurately reconstructs the low-frequency target, with $\mathcal{S}(k)$ progressively increasing in the relevant modes. However, for the high-frequency target, the output remains near zero and $\mathcal{S}(k)$ is steadily suppressed, indicating an inability to fit high-frequency components. To address this phenomenon, we introduce a weight perturbation scheme in section IV.
@@
Remark. As observed in Figs. 3 and 4, an extreme case of a spectral bottleneck is characterized by: (a) progressive attenuation of the NTK eigenmode energy with each training epoch, (b) failure of the network to represent even the low-frequency components (highlighted in Fig. 4) that lie within the nominal frequency support of SIREN, and (c) convergence to a zero-valued output.
IV WINNER: Weight Initialization with Noise for NEural Representations